# Happy family of stable marriages

### 02 Apr 2026 v1

@@ 6.1 Background and Motivation
In classical convex geometry, the Minkowski and Alexandrov problems seek to reconstruct a convex body from its standard Euclidean surface area measure or integral Gauss curvature, respectively. In this framework, we introduce a novel variation: an Alexandrov-type problem formulated with respect to the self-surface measure of a centrally symmetric convex set.
Instead of relying on the Euclidean surface area, the perimeter and volume are defined intrinsically. For a centrally symmetric convex set $K\subset\mathbb{R}^{n}$, the self-perimeter is constructed recursively by integrating over the essential boundary, normalized by the self-volume of the (n1)(n-1)-dimensional cross-sections. The inverse problem therefore asks: Given a probability measure $\mu$ on $S^{n-1}$, does there exist a constant $\lambda>0$ and a centrally symmetric convex set $B\subset\mathbb{R}^{n}$ whose self-surface measure is exactly $\lambda\mu$?
@@
Let $\mu$ be a measure on $S^{n-1}$. For a centrally symmetric convex set $K$ let $T_{B}:S^{n-1}\rightarrow\partial B$ be given by $T_{B}(\theta)=r_{K}(\theta)\theta$ (see Definition 2.1). Let $\nu_{B}(\theta)$ be the exterior normal to $\partial B$ at the point $x=T_{B}(\theta)$ . Using the Jacobian of of the radial map $d{\cal H}_{n-1}(x)=\frac{r_{B}^{n-1}(\theta)}{<\theta,\nu_{B}(\theta)>}d\theta$ we may convert (11) to
$P_{B}^{(n)}:=\int_{S^{(n-1)}}\frac{V(B^{(n-1)}(\nu_{B}(\theta)))}{{\cal H}_{n-1}(B^{(n-1)}(\nu_{B}(\theta))))}\frac{r_{B}^{n-1}(\theta)}{<\theta,\nu_{B}(\theta)>}d\theta\ .$
Alexandrov’s type problem:
Let $\mu$ be a positive measure on $S^{n-1}$. Find a central symmetric convex set $B$ in $\mathbb{R}^{n}$ whose self surface measure is consistent with $\mu$, namely
$\mu(d\theta)=\frac{V(B^{(n-1)}(\nu_{B}(\theta)))}{{\cal H}_{n-1}(B^{(n-1)}(\nu_{B}(\theta))))}\frac{r_{B}^{n-1}(\theta)}{<\theta,\nu_{B}(\theta)>}d\theta\ .$ (24)
Using (3) we pose the Alexandrov type problem in $\mathbb{R}^{2}$: Let $\mu$ be a $\pi$ periodic density on $S^{1}$ . Is there a central convex set $B$ such that
$\mu(d\theta)=\frac{\sqrt{r^{2}_{B}(\theta)+(r^{{}^{\prime}}_{B}(\theta))^{2}}d\theta}{r_{B}\left(\theta+\arctan\left(\frac{r_{B}(\theta}{r_{B}^{{}^{\prime}}(\theta)})\right)\right)}\ \$(25)
Let $\mu:=e^{\epsilon(\phi+\phi_{0})}d\theta$ where $\phi_{0}\in\mathbb{R}$ and $\int_{0}^{2\pi}\phi d\theta=0$. We may write $\phi=\phi^{(p)}+\phi^{(u)}$ where
$\phi^{(p)}=\sum_{k\in 4\mathbb{Z}-0}\phi_{k}e^{ik\theta}\ ;\ \ \phi^{(u)}=\sum_{k\not\in 4\mathbb{Z}}\phi_{k}e^{ik\theta}\ \ ;\phi_{k}=\phi_{-k}^{*}$
@@
There exists a unique $\phi_{0}\in\mathbb{R}$ such that
@@
$\Gamma(\gamma):=\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)-\gamma>0\}}\sqrt{\phi^{(p)}(s)-\gamma}\,ds-\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)-\gamma<0\}}\sqrt{-\phi^{(p)}(s)+\gamma}\,ds$
It follows the $\Gamma$ is continuous, monotone decreasing, $\Gamma$ is negative for $\gamma>>1$ and positive for $\gamma<<1$, so the lemma is obtained by mean value theorem. ∎
@@
If $\phi^{(p)}\not\equiv 0$ and $\phi_{0}$ satisfies (26) then the asymptotic expansion of the Alexandrov problem around the unit circle $r_{B}\equiv 1$ is
@@
Let $\zeta=\ln(r_{B})$ where $r_{B}$ a solution of (26).
Taking the log of both sides we get
@@
Let $L(\zeta)_{(\theta)}:=\zeta(\theta)-\zeta(\theta+\pi/2)$. The kernel of $L$ is the set of $\pi/2$ periodic functions. Its spectrum is $\lambda_{k}=(1-e^{ik\pi/2})$ and the eigenfunctions are just $e^{ik\theta}$, $k\in\mathbb{Z}$.
To balance the equation, we perform an asymptotic expansion $\zeta=\epsilon^{1/2}\zeta_{1}+\epsilon\zeta_{2}+o(\epsilon)$.
At $\mathcal{O}(\epsilon^{1/2})$, $L\zeta_{1}=0$. This implies $\zeta_{1}$ is purely $\pi/2$-periodic, containing only harmonics where $k\in 4\mathbb{Z}$. It implies $\zeta_{1}^{{}^{\prime}}(\theta+\pi/2)=\zeta_{1}^{{}^{\prime}}(\theta)$ so by (28)
@@
The right side of (29) should contain only the harmonics with $k\not\in 4\mathbb{Z}$. Thus
$(\zeta_{1}^{{}^{\prime}})^{2}=\frac{2}{3}(\phi^{(p)}+\phi_{0})\ \rightarrow\zeta_{1}^{{}^{\prime}}=\pm sign(\phi^{(p)}+\phi_{0})\sqrt{\frac{2}{3}|\phi^{(p)}+\phi_{0}|}$
Since $\zeta_{1}$ is $2\pi$ periodic it implies $\int_{0}^{2\pi}\zeta_{1}^{{}^{\prime}}(\theta)d\theta=0$. This is satisfied iff $\phi_{0}$ satisfies (26). ∎
@@
By (29) we also get
$\zeta_{2}(\theta)=\sum_{k\not\in 4\mathbb{Z}}\frac{\phi_{k}e^{ik\theta}}{1-e^{ik\pi/2}}\mod Ker(L)$
@@ 7 Conclusion
In this paper, we introduced a natural, recursive definition for volume and surface area in $n$-dimensional normed spaces, centered on preserving the Euclidean ratio $P(B)/V(B)=n$ We demonstrated that this construction is invariant under origin-preserving affine transformations and polar duality, satisfying fundamental geometric requirements for measures in Minkowski spaces [1, 5].
For the two-dimensional case, we utilized the property that the self-perimeter is a strictly convex functional with respect to the interior reference point, which ensures the existence of a unique minimizing generalized centroid for any convex body [8]. We further discussed the established result that the minimum self-perimeter of any planar convex body is bounded above by 9, with the maximum value attained if and only if the body is a triangle evaluated at its centroid [7]. This sharp upper bound complements the classical results for centrally symmetric sets [6].
Finally, we explored an Alexandrov-type problem for the associated self-surface measure [4]. Through a perturbative analysis around the Euclidean disc, we identified a rigidity result for the surface measure $\mu$:
significant obstruction to solvability: a necessary condition for the existence of a solution is the absence of certain harmonic components in the density perturbation. These findings open new avenues for research, particularly regarding the global solvability of the inverse problem and the extremal values of the self-volume in higher dimensions.

### 09 Apr 2026 v2

@@ 5.1 Background and Motivation
In classical convex geometry, the Minkowski and Alexandrov problems seek to reconstruct a convex body from its standard Euclidean surface area measure or integral Gauss curvature, respectively. In this framework, we introduce a novel variation: an Alexandrov-type problem formulated with respect to the self-surface measure of a centrally symmetric convex set. Instead of relying on the Euclidean surface area, the perimeter and volume are defined intrinsically. For a centrally symmetric convex set $K\subset\mathbb{R}^{n}$, the self-perimeter is constructed recursively by integrating over the essential boundary, normalized by the self-volume of the (n1)(n-1)-dimensional cross-sections. The inverse problem therefore asks: Given a measure $\mu$ on $S^{n-1}$, does there exist a centrally symmetric convex set $B\subset\mathbb{R}^{n}$ whose self-surface measure is exactly $\lambda\mu$?
@@
Let $\mu$ be a measure on $S^{n-1}$. For a centrally symmetric convex set $B$, let $T_{B}:S^{n-1}\rightarrow\partial B$ be given by $T_{B}(\theta)=r_{B}(\theta)\theta$ (see Definition 2.1). Let $\nu_{B}(\theta)$ be the exterior normal to $\partial B$ at the point $x=T_{B}(\theta)$. Using the Jacobian of the radial map $d{\cal H}_{n-1}(x)=\frac{r_{B}^{n-1}(\theta)}{<\theta,\nu_{B}(\theta)>}d\theta$, we may convert (13) to:
$\omega^{(n)}_{B}:=n^{-1}\int_{S^{(n-1)}}\frac{\omega^{(n-1)}(B^{(n-1)}(\nu_{B}(\theta)))}{{\cal H}_{n-1}(B^{(n-1)}(\nu_{B}(\theta))))}\frac{r_{B}^{n-1}(\theta)}{<\theta,\nu_{B}(\theta)>}d\theta\ .$
Alexandrov-type problem: Let $\mu$ be a positive measure on $S^{n-1}$. Find a centrally symmetric convex set $B$ in $\mathbb{R}^{n}$ whose self-surface measure is consistent with $\mu$, namely:
$\mu(d\theta)=\frac{\omega^{(n-1)}(B^{(n-1)}(\nu_{B}(\theta)))}{{\cal H}_{n-1}(B^{(n-1)}(\nu_{B}(\theta))))}\frac{r_{B}^{n-1}(\theta)}{<\theta,\nu_{B}(\theta)>}d\theta\ .$ (18)
Using (4), we pose the Alexandrov-type problem in $\mathbb{R}^{2}$: Let $\mu$ be a $\pi$-periodic density on $S^{1}$. Is there a central convex set $B$ such that:
$\mu(d\theta)=\frac{\sqrt{r^{2}_{B}(\theta)+(r^{{}^{\prime}}_{B}(\theta))^{2}}d\theta}{r_{B}\left(\theta+\arctan\left(\frac{r_{B}(\theta)}{r_{B}^{{}^{\prime}}(\theta)}\right)\right)}\ \$(19)
Let $\mu:=e^{\epsilon(\phi+\phi_{0})}d\theta$ where $\phi_{0}\in\mathbb{R}$ and $\int_{0}^{2\pi}\phi d\theta=0$. We may write $\phi=\phi^{(p)}+\phi^{(u)}$ where:
$\phi^{(p)}=\sum_{k\in 4\mathbb{Z}-\{0\}}\phi_{k}e^{ik\theta}\ ;\ \ \phi^{(u)}=\sum_{k\not\in 4\mathbb{Z}}\phi_{k}e^{ik\theta}\ \ ;\phi_{k}=\phi_{-k}^{*}$
@@
There exists a unique $\phi_{0}\in\mathbb{R}$ such that:
@@
$\Gamma(\gamma):=\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)-\gamma>0\}}\sqrt{\phi^{(p)}(s)-\gamma}\,ds-\int_{\{s\in\mathbb{S}^{1},\phi^{(p)}(s)-\gamma<0\}}\sqrt{-(\phi^{(p)}(s)-\gamma)}\,ds$
It follows that $\Gamma$ is continuous, monotone decreasing, $\Gamma$ is negative for $\gamma>>1$ and positive for $\gamma<<1$, so the lemma is obtained by the mean value theorem. ∎
@@
If $\phi^{(p)}\not\equiv 0$ and $\phi_{0}$ satisfies (20), then the asymptotic expansion of the Alexandrov problem around the unit circle $r_{B}\equiv 1$ is:
@@
Let $\zeta=\ln(r_{B})$ where $r_{B}$ is a solution. Taking the log of both sides we get:
@@
Let $L(\zeta)_{(\theta)}:=\zeta(\theta)-\zeta(\theta+\pi/2)$. The kernel of $L$ is the set of $\pi/2$-periodic functions. Its spectrum is $\lambda_{k}=(1-e^{ik\pi/2})$ and the eigenfunctions are just $e^{ik\theta}$, $k\in\mathbb{Z}$. To balance the equation, we perform an asymptotic expansion $\zeta=\epsilon^{1/2}\zeta_{1}+\epsilon\zeta_{2}+o(\epsilon)$. At $\mathcal{O}(\epsilon^{1/2})$, $L\zeta_{1}=0$. This implies $\zeta_{1}$ is purely $\pi/2$-periodic, containing only harmonics where $k\in 4\mathbb{Z}$. It implies $\zeta_{1}^{{}^{\prime}}(\theta+\pi/2)=\zeta_{1}^{{}^{\prime}}(\theta)$ so by (22):
@@
The right side of (23) should contain only the harmonics with $k\not\in 4\mathbb{Z}$. Thus:
$(\zeta_{1}^{{}^{\prime}})^{2}=\frac{2}{3}(\phi^{(p)}+\phi_{0})\ \rightarrow\zeta_{1}^{{}^{\prime}}=\pm\operatorname{sgn}(\phi^{(p)}+\phi_{0})\sqrt{\frac{2}{3}|\phi^{(p)}+\phi_{0}|}$
Since $\zeta_{1}$ is $2\pi$-periodic it implies $\int_{0}^{2\pi}\zeta_{1}^{{}^{\prime}}(\theta)d\theta=0$. This is satisfied iff $\phi_{0}$ satisfies (20). ∎
@@
By (23) we also get:
$\zeta_{2}(\theta)=\sum_{k\not\in 4\mathbb{Z}}\frac{\phi_{k}e^{ik\theta}}{1-e^{ik\pi/2}}\mod\text{Ker}(L)$
@@ 6 Conclusion
In this paper, we introduced a natural, recursive definition for volume and surface area in $n$-dimensional normed spaces, centered on preserving the Euclidean ratio $P(B)/\omega^{(n)}_{B}(B)=n$ [2, 3]. We demonstrated that this construction is invariant under origin-preserving affine transformations and polar duality, satisfying fundamental geometric requirements for measures in Minkowski spaces [1, 5]. For the two-dimensional case, we utilized the property that the self-perimeter is a strictly convex functional with respect to the interior reference point, which ensures the existence of a unique minimizing generalized centroid for any convex body [8]. We further discussed the established result that the minimum self-perimeter of any planar convex body is bounded above by 9, with the maximum value attained if and only if the body is a triangle evaluated at its centroid [7]. This sharp upper bound complements the classical results for centrally symmetric sets [6]. Finally, we explored an Alexandrov-type problem for the associated self-surface measure [4]. Through a perturbative analysis around the Euclidean disc, we identified a rigidity result for the surface measure $\mu$, revealing a significant obstruction to solvability: a necessary condition for the existence of a solution is the absence of certain harmonic components in the density perturbation. These findings open new avenues for research, particularly regarding the global solvability of the inverse problem and the extremal values of the self-volume in higher dimensions.
Acknowledgments
The author wishes to acknowledge Jenni (also known as Gemini) for her invaluable assistance in editing, proofreading, and structuring this manuscript.